An algorithm takes 0.5 ms for input size of 100. How long will it take for an input of size of 500 assuming a big-O runtime of:

a. O(? 2 )

b. O(? log10 ?)
Answer #1

a)

t=\frac{n_1^2}{n_2^2}*t_2

t= 5002 * 0.5 = 25 * 0.5 = 12.5 ms

Answer: 12.5 ms

b)

nalog 10 11 nalog 10 n2 to

500 log 10500 - *0.5 100 log 10 100

t = 5*1.35 * 0.5 = 3.375ms

Answer: 3.375 ms